# Explicit vs. Implicit: Investigating Social Bias in Large Language Models through Self-Reflection

### v2

@@ Abstract
Large Language Models (LLMs) have been shown to exhibit various biases and stereotypes in their generated content. While extensive research has investigated bias in LLMs, prior work has predominantly focused on explicit bias, leaving the more nuanced implicit biases largely unexplored. This paper presents a systematic framework grounded in social psychology theories to investigate and compare explicit and implicit biases in LLMs. We propose a novel "self-reflection" based evaluation framework that operates in two phases: first measuring implicit bias through simulated psychological assessment methods, then evaluating explicit bias by prompting LLMs to analyze their own generated content. Through extensive experiments on state-of-the-art LLMs across multiple social dimensions, we demonstrate that LLMs exhibit a substantial inconsistency between explicit and implicit biases, where explicit biases manifest as mild stereotypes while implicit biases show strong stereotypes. Furthermore, we investigate the underlying factors contributing to this explicit-implicit bias inconsistency. Our experiments examine the effects of training data scale, model parameters, and alignment techniques. Results indicate that while explicit bias diminishes with increased training data and model size, implicit bias exhibits a contrasting upward trend. Notably, contemporary alignment methods (e.g., RLHF, DPO) effectively suppress explicit bias but show limited efficacy in mitigating implicit bias. These findings suggest that while scaling up models and alignment training can address explicit bias, the challenge of implicit bias requires novel approaches beyond current methodologies.
@@
Yachao Zhao, Bo Wangthanks: *Corresponding author., Yan Wang, College of Intelligence and Computing, Tianjin University, Tianjin, China {zhaoyachao, bo_wang}@tju.edu.cn
@@ 1 Introduction
Figure 1: Our proposed assessment methodology based on self-reflection. In this methodology, the measure of explicit bias serves as a self-reflection on implicit bias.
@@
Research in social bias within Large Language Models (LLMs) has made substantial progress, with researchers developing various methods to identify and quantify model biases. However, current bias research faces significant limitations - these studies either fail to differentiate between explicit and implicit biases in LLMs, or predominantly focus on explicit bias while overlooking implicit bias. These limitations result in an incomplete understanding of model bias, highlighting the need for a more comprehensive research on the relation between explicit and implicit bias.
@@
In social psychology, social bias is categorized into two forms, explicit and implicit biasGreenwald and Banaji (1995). Explicit bias, typically manifesting as conscious and controllable attitude expressions, is primarily measured through Self-Report Assessment (SRA) Lajunen and Özkan (2011). In contrast, implicit bias, which reflects unconscious and automated cognitive processes, is commonly measured by the Implicit Association Test (IAT) Greenwald et al. (1998). Extensive psychological research examining the relation between explicit and implicit biases in human individuals has revealed frequent inconsistencies between these two forms of bias, particularly regarding sensitive social targets such as gender and race. Furthermore, psychological research has extensively investigated the causes of this inconsistency, primarily attributing it to early learning processesBaron and Banaji (2006) and social expectationsCrandall et al. (2002). However, these investigations remain predominantly focused on human individuals, with limited research examining the explicit and implicit biases exhibited by large language models.
Based on these theoretical foundations, this study proposes an evaluation methodology based on prompt templates and self-reflection. This approach maps the measurement of explicit and implicit biases in LLMs to psychological assessment methods: SRA and IAT, respectively, thereby extending evaluation methods from human subjects to LLMs. The core of this methodology lies in measuring explicit bias through model self-reflection (Self-reflection) on its manifested implicit biases. It enables the comparative analysis of explicit versus implicit biases regarding identical targets.
@@
To comprehensively evaluate and compare explicit and implicit biases in LLMs, this study conducted experiments on a representative selection of models, including both proprietary and open-source mainstream LLMs such as ChatGPTOuyang et al. (2022), Claude-3.5-SonnetBai et al. (2022b), and LLaMA-3.1-405B-InstructDubey et al. (2024). The experiments encompassed multiple social dimensions including gender, race, occupation, age, and disability. Results reveal a phenomenon that all LLMs exhibit bias inconsistency patterns similar to humans across different dimensions: demonstrating minimal stereotyping at the explicit level while manifesting significant stereotypical associations at the implicit level.
@@
Given these findings, this study further investigates the underlying causes of bias inconsistency in LLMs. While social psychology has extensively explored similar inconsistency in human individuals, the mechanisms behind such patterns in LLMs remain unclear. This study systematically examines three key factors that potentially contribute to bias inconsistency in LLMs: training data scale, model size, and alignment training. Our investigation focuses on the LLaMA model familyTouvron et al. (2023); Dubey et al. (2024), which provides an ideal experimental setting due to its open-source nature, transparent training data documentation, and availability of multiple model sizes with both base and instruction-tuned versions. Our experimental results reveal complex relation between these factors and bias manifestation in LLMs. As training data scale and model parameters increase, explicit bias shows a consistent decrease, while implicit bias demonstrates the opposite trend, showing steady increase. Alignment training exhibits a distinct impact - while it significantly reduces explicit bias, implicit bias remains relatively stable within a fixed range regardless of training steps. These findings suggest that while recent advances in LLM development have successfully reduced explicit biases through scaling and alignment techniques, addressing implicit biases may require fundamentally different approaches.
In summary, the main contributions of this paper are as follows:
@@
1. We propose a novel "self-reflection" evaluation methodology aligned with psychological theory. Through carefully designed prompt templates, we employ SRA to measure explicit bias and IAT to measure implicit bias. Moreover, the assessment of explicit bias involves the model’s self-reflection on its implicit biases.
@@
2. Our experimental results reveal inconsistency between explicit and implicit biases in LLMs. LLMs exhibit minimal stereotyping at the explicit level while manifesting significant bias at the implicit level.
@@
3. We present the first systematic investigation of key factors influencing explicit and implicit biases in LLMs. Through in-depth experimental analysis across three dimensions - training data scale, model size, and alignment training - we uncover the differential impact mechanisms these factors have on both types of bias.
@@ 2.1 Bias Research in Social Psychology
In social psychology, research on attitudinal bias represents a crucial area of investigation. Bias is commonly defined as preferential views toward specific targets Smith and Noble (2014). Researchers distinguish between two forms of bias: explicit and implicit Greenwald and Banaji (1995). Explicit bias refers to consciously recognized and deliberately expressed biased attitudes, which are overt and conscious in nature, primarily measured through Self-Report Assessment (SRA) Lajunen and Özkan (2011). The Likert Scale Likert (1932) is one of the most widely used measurement instruments in SRA, evaluating participants’ attitudes through 5-point or 7-point rating scales. Implicit bias refers to attitudes that individuals find difficult to identify and are typically unaware of, manifesting through underlying associations or automatic responses. These biases are commonly measured using the Implicit Association Test (IAT) Greenwald et al. (1998). In a typical IAT experiment, researchers evaluate implicit attitudes by measuring participants’ response times to different concept pairings. Social psychologists have conducted extensive comparative studies on explicit and implicit biases in individuals Dovidio et al. (2001); Nosek (2007); Son Hing et al. (2008). These studies have revealed a significant finding: when addressing sensitive social topics such as gender and race, individuals often demonstrate notable discrepancies between their explicit and implicit biases. For instance, research has shown that while participants explicitly express support for gender equality, their implicit attitudes reveal negative perceptions toward female scientists Moss-Racusin et al. (2012). This divergence is primarily attributed to individuals’ socialization processes and social expectations. Implicit biases typically form during early childhood, while the development of cognitive abilities enables individuals to progressively suppress the explicit expression of these implicit biases, resulting in this inconsistencyBaron and Banaji (2006). Additionally, this inconsistency is further amplified when mainstream social values conflict with implicit individual attitudesCrandall et al. (2002).
@@ 2.2 Bias Analysis in Large Language Models
Prior work has extensively investigated bias in Large Language Models. Some research has explored explicit bias in LLMs, typically measuring bias by directly including specific target objects in prompts or templatesAbid et al. (2021); Kirk et al. (2021); Kotek et al. (2023). Abid et al. Abid et al. (2021) found that GPT-3 Brown et al. (2020) tends to generate text with violence when given Muslim-related prompts. Other studies examine models’ implicit biasesCaliskan et al. (2017); Zhao et al. (2024). Marked PersonasCheng et al. (2023), through analyzing distributional differences in model-generated words across social groups, revealed that apparently positive model outputs contained implicit negative perceptions of Black women. Bai et al.Bai et al. (2024) developed a methodology based on the Implicit Association Test (IAT) to quantify implicit biases embedded in LLMs.
Beyond these bias detection methods, studies have also focused on investigating the factors influencing social bias in LLMs. Inspired by the Scaling Laws Kaplan et al. (2020), these studies primarily examine the relation between bias and various factors such as model size and training data composition. Furthermore, with the widespread use of Reinforcement Learning from Human Feedback (RLHF) Ouyang et al. (2022) in LLM training, investigations have emerged to understand how alignment methods affect the bias in LLMs. Ganguli et al. Ganguli et al. (2023) focused on two key variables: model scale and RLHF training steps, examining their impact on models’ Moral Self-Correction capabilities. Their large-scale experiments revealed that within certain ranges, both model scale and RLHF training steps showed significant positive correlations with models’ ability to self-debias. Through a series of experiments, Ali et al. Ali et al. (2024) found that increasing model size may lead to higher levels of bias.
@@ 3 Methodology
Large language models(LLMs) possess the capability to reflect on and evaluate their own outputs, providing a new perspective for studying their biasesMadaan et al. (2023); Weng et al. (2023); Shinn et al. (2023). The evaluation method proposed in this study fully utilizes the model’s reflective capabilities: first measuring the model’s implicit biases, then measuring its explicit biases through the model’s reflection on these implicit biases. This method effectively compares both implicit and explicit biases toward the same social targets within the model, revealing the relation between these two type of biases.
Specifically, the evaluation method introduced in this section draws inspiration from psychological measures including the Implicit Association Test (IAT) and Self-Report Assessment (SRA), which are used to measure implicit and explicit biases respectively. Through carefully designed prompt engineering, these methods are adapted for evaluating LLMs.
@@ 3.1 Implicit Bias Assessment Based on Psychological Theory
Research in social psychology suggests that the key to measuring implicit bias lies in conducting assessments without subjects’ awareness Greenwald and Banaji (1995). The Implicit Association Test (IAT) is an important measurement tool designed based on this principle: rather than directly inquiring about subjects’ attitudes toward social groups such as gender or race, it measures implicit bias by examining differences in response times when subjects associate target-related stimuli (such as names of different genders/races) with corresponding attributes. For example, if a subject’s response time is significantly shorter for "male-career, female-family" combinations compared to "female-career, male-family" combinations, this may indicate an implicit bias associating men with careers and women with family roles.
To measure implicit bias in LLMs, this study innovatively adapts the IAT to model evaluation through prompt engineering. Specifically, we propose a prompt template containing attribute words and masked targets:
<mask> is attrX as <mask> is attrY,   (1) where $mask$ represents the masked target, while $attr_{X}$ and $attr_{Y}$ represent a predefined pair of attributes. This template requires the model to make a choice among given candidate words, thereby revealing implicit bias. To illustrate the methodology, we examine implicit associations between racial categories (White-Black) and valence attributes (good-bad). The experimental procedure consists of the following steps:
@@
    Define attribute word pairs $A_{X}$ and $A_{Y}$, representing ’good’ and ’bad’ respectively.
@@
    Randomly select N stimuli from $S_{a}$ and $S_{b}$ to form subsets $s_{a}$ and $s_{b}$, and randomly choose one attribute word each ($a_{X}$ and $a_{Y}$) from $A_{X}$ and $A_{Y}$.
@@
    Insert $a_{X}$ and $a_{Y}$ into the attribute positions of the template, and combine and shuffle $s_{a}$ and $s_{b}$ to create a candidate word set $s_{ab}$.
@@
    Design a prompt including the above elements with corresponding instructions, requiring the model to select two words from $s_{ab}$ to fill the $mask$ positions in the template.
  6. 6.
    To enhance measurement reliability, following recommendations on prompt template diversification from existing research Seshadri et al. (2022); Goldfarb-Tarrant et al. (2023), we implemented several optimization methods: First, we created 5 semantically similar but structurally distinct test templates through synonym substitution and word modification. To further mitigate order effects, we swapped the attribute pairs in these 5 templates, resulting in 10 templates.
Beyond racial bias, this study conducts comparative experiments on both explicit and implicit biases across multiple sensitive social dimensions including gender, occupation, age, and disability, ensuring comprehensive experimental coverage.
@@ 3.2 Explicit Bias Assessment Through Self-Reflection
To measure explicit bias in LLM, this study draws on Self-report Assessment (SRA) from social psychology, a classic technique for measuring explicit bias. Unlike the indirect measurement of implicit bias, explicit bias measurement requires subjects to directly express their attitudes and views toward specific social groups while being fully aware of the measurement purpose. The Likert Scale Likert (1932) is one of the most commonly used measurement tools in self-report assessment, evaluating subjects’ attitudes using 5-point or 7-point rating scales.
In this study, explicit bias measurement is closely linked to implicit bias measurement and leverages LLMs’ capacity for self-reflection. Specifically, this paper measures explicit bias by guiding language models to systematically reflect on and evaluate their potential attitudes demonstrated during implicit bias measurement. Because this process requires LLMs to analyze and evaluate their implicit biases shown in Section 3.1, we term this approach the "self-reflection" assessment method.
In the implementation process, the measurement template needs to explicitly specify the target social groups, thus requiring the replacement of $mask$ in template (1) from the implicit test with specific social group concept words, such as "men" and "women" for gender categories, or "Black" and "White" for racial categories. Specifically, this paper randomly selects words from the predefined social group word sets $T_{A}$ and $T_{B}$ to replace $mask$, while keeping the predefined attribute words $attribute_{X}$ and $attribute_{Y}$ unchanged, resulting in the following template sentence:
TargetA is attrX as TargetB is attrY,    (2)
Subsequently, the LLM is prompted to evaluate the template sentence based on a five-point Likert scale, with options including: strongly disagree, disagree, neutral, agree, strongly agree. The model should assess its level of agreement based on the specific content of the sentence, select the most appropriate option from these five choices, and provide reasoning for its selection.
To enhance measurement reliability and mitigate potential order effects, this study randomizes the five Likert scale options during each test rather than using a fixed order. This operation helps reduce potential interference from option ordering on model choices, thereby maximizing the stability and consistency of measurement results.

### v4

@@ Abstract
Large Language Models (LLMs) have been shown to exhibit various biases and stereotypes in their generated content. While extensive research has investigated biases in LLMs, prior work has predominantly focused on explicit bias, with minimal attention to implicit bias and the relation between these two forms of bias. This paper presents a systematic framework grounded in social psychology theories to investigate and compare explicit and implicit biases in LLMs. We propose a novel self-reflection-based evaluation framework that operates in two phases: first measuring implicit bias through simulated psychological assessment methods, then evaluating explicit bias by prompting LLMs to analyze their own generated content. Through extensive experiments on advanced LLMs across multiple social dimensions, we demonstrate that LLMs exhibit a substantial inconsistency between explicit and implicit biases: while explicit bias manifests as mild stereotypes, implicit bias exhibits strong stereotypes. We further investigate the underlying factors contributing to this explicit-implicit bias inconsistency, examining the effects of training data scale, model size, and alignment techniques. Experimental results indicate that while explicit bias declines with increased training data and model size, implicit bias exhibits a contrasting upward trend. Moreover, contemporary alignment methods effectively suppress explicit bias but show limited efficacy in mitigating implicit bias.
@@
Yachao Zhao1, Bo Wang1thanks: Corresponding author., Yan Wang1, Dongming Zhao2, Ruifang He1, Yuexian Hou1 1College of Intelligence and Computing, Tianjin University, Tianjin, China 2AI Lab, China Mobile Communication Group Tianjin Co., Ltd. {zhaoyachao, bo_wang}@tju.edu.cn
@@ 1 Introduction
Figure 1: Our proposed assessment methodology based on psychological theories and self-reflection. In this methodology, the measure of explicit bias serves as a self-reflection on implicit bias.
@@
Research on social bias within Large Language Models (LLMs) has made substantial progress, with a variety of methods now available to identify and quantify biases. However, these works face significant limitations: they either fail to differentiate between explicit and implicit biases in LLMs or predominantly focus on explicit bias while overlooking implicit bias and the relation between the two types of bias. These limitations result in an incomplete understanding of model biases, highlighting the need for more comprehensive research that explores both types of bias and the connections between them.
@@
In social psychology, social bias is categorized into two forms, explicit and implicit bias Greenwald and Banaji (1995). Explicit bias, typically manifesting as conscious attitudinal tendencies, is primarily measured through Self-Report Assessment (SRA) Lajunen and Özkan (2011). In contrast, implicit bias, which reflects unconscious and automated cognitive processes, is commonly measured by the Implicit Association Test (IAT) Greenwald et al. (1998). Numerous psychological studies investigate the relation between explicit and implicit biases in human individuals. They find frequent inconsistencies between these two forms of bias, particularly regarding sensitive social targets such as gender and race. Researchers have further explored the causes of these inconsistencies, primarily attributing them to early learning processes Baron and Banaji (2006) and social expectations Crandall et al. (2002). However, these investigations remain predominantly focused on human individuals, with limited research studying the explicit and implicit biases exhibited by LLMs.
@@
Drawing from these theoretical foundations, we propose an evaluation methodology for LLMs that leverages prompt templates and self-reflection. This approach maps the measurement of explicit and implicit biases within LLMs to psychological assessment methods: SRA and IAT, thereby extending evaluation methods from human subjects to LLMs. The core of this methodology lies in measuring explicit bias through LLMs’ self-reflection on their expressed implicit bias. It enables a comparative analysis of explicit and implicit biases toward identical targets. To comprehensively evaluate and compare explicit and implicit biases in LLMs, we conduct experiments on a representative selection of models, including both proprietary and open-source mainstream LLMs such as ChatGPT Ouyang et al. (2022), Claude-3.5-Sonnet Bai et al. (2022b), and LLaMA-3.1-405B-Instruct Dubey et al. (2024). Our experiments encompass multiple social dimensions including gender, race, occupation, age, and disability. The results reveal that all tested LLMs exhibit bias inconsistency patterns similar to those observed in humans across different dimensions: demonstrating minimal stereotyping at the explicit level while manifesting significant stereotypical associations at the implicit level.
@@
Given these findings, we further investigate the underlying causes of bias inconsistency in LLMs. Although social psychology has widely explored similar inconsistencies in human individuals, the mechanisms driving such patterns in LLMs remain unclear. To address this gap, we systematically examine three key factors that potentially contribute to bias inconsistency in LLMs: training data scale, model size, and alignment training. Our study focuses on the LLaMA model family Touvron et al. (2023); Dubey et al. (2024), which provides an ideal experimental setting due to its open-source nature, transparent training data documentation, and availability of multiple model sizes with both base and instruction-tuned versions. Experimental results reveal a complex relation between these factors and bias manifestation in LLMs. Specifically, as training data scale and model size increase, explicit bias shows a consistent decrease while implicit bias demonstrates a steady increase. Alignment training, on the other hand, has a distinctive impact: it significantly reduces explicit bias, while implicit bias remains relatively stable within a fixed range regardless of training steps. Under the combined influence of these factors, LLMs display an inconsistency between their explicit and implicit biases. These findings suggest that although recent advancements in LLMs have successfully reduced explicit bias, addressing implicit bias may still require fundamentally different approaches. In summary, the main contributions of this paper are as follows:
@@
1. We propose a novel self-reflection-based evaluation methodology aligned with psychological theories. Through prompt templates, we employ SRA to measure explicit bias and IAT to measure implicit bias. Moreover, our assessment of explicit bias involves the LLM’s self-reflection on its implicit bias.
@@
2. Our experimental results reveal an inconsistency between explicit and implicit biases in LLMs. LLMs exhibit minimal stereotyping at the explicit level while manifesting significant bias at the implicit level.
@@
3. We present the first systematic investigation of key factors influencing explicit and implicit biases in LLMs. Through in-depth experimental analysis across three dimensions: training data scale, model size, and alignment training, we uncover how these factors impact both types of bias and drive the inconsistency between them.
@@ 2.1 Bias Research in Social Psychology
In social psychology, research on attitudinal bias represents a crucial area of investigation. Bias is commonly defined as preferential views toward specific targets Smith and Noble (2014), and researchers distinguish between two primary forms: explicit and implicit bias Greenwald and Banaji (1995). Explicit bias refers to consciously recognized and deliberately expressed attitudes of preference. It is primarily measured through Self-Report Assessment (SRA) Lajunen and Özkan (2011), with the Likert scale Likert (1932) serving as one of the most prevalent measurement instruments, employing 5-point or 7-point rating scales to evaluate participants’ attitudes. In contrast, implicit bias refers to attitudes of preference that individuals struggle to identify and are typically unaware of, manifesting through underlying associations or automatic responses. It is commonly measured using the Implicit Association Test (IAT) Greenwald et al. (1998), which evaluates implicit attitudes by analyzing participants’ response times to different concept pairings.
Social psychologists have conducted extensive comparative studies on explicit and implicit biases Dovidio et al. (2001); Nosek (2007); Son Hing et al. (2008). These studies reveal a significant finding: when addressing sensitive social topics such as gender and race, individuals often demonstrate notable discrepancy between their explicit and implicit biases. For instance, while participants may explicitly express support for gender equality, their implicit attitudes betray negative perceptions toward female scientists Moss-Racusin et al. (2012). Researchers have further explored the causes of this divergence, attributing it primarily to socialization processes and social expectations. Specifically, implicit biases typically form during early childhood, whereas the development of cognitive abilities enables individuals to progressively suppress the explicit expression of these biases, thereby resulting in this inconsistency Baron and Banaji (2006). Additionally, this inconsistency is further amplified when mainstream social values conflict with implicit individual attitudes Crandall et al. (2002).
@@ 2.2 Bias Analysis in Large Language Models
Prior works have widely investigated bias in Large Language Models (LLMs). Some researches explore explicit bias, typically measuring it by directly including specific target objects in prompts or templates Abid et al. (2021); Kirk et al. (2021); Kotek et al. (2023). For instance, Abid et al. (2021) found that GPT-3 Brown et al. (2020) tends to generate text with violence when given Muslim-related prompts. Other studies examine models’ implicit biases Caliskan et al. (2017); Cheng et al. (2023); Zhao et al. (2024). Bai et al. (2024) developed a methodology based on the Implicit Association Test (IAT) to quantify implicit biases embedded in LLMs. Beyond these detection methods, there have been investigations into factors influencing social bias in LLMs. Inspired by the Scaling Laws Kaplan et al. (2020), these studies primarily center on the relation between bias and various factors such as model size and training data composition. Furthermore, with the widespread use of Reinforcement Learning from Human Feedback (RLHF) Ouyang et al. (2022), new works have emerged to understand how alignment methods affect the bias in LLMs. Ganguli et al. (2023) focused on two key variables: model scale and RLHF training steps, analyzing their impact on models’ moral self-correction capabilities. Their large-scale experiments reveal that within certain ranges, both model scale and RLHF training steps showed significant positive correlations with models’ ability to self-debias. Through a series of experiments, Ali et al. (2024) found that increasing model size may lead to more severe bias.
Despite significant progress in bias detection in LLMs, existing works often examine explicit bias and implicit bias as independent research subjects, lacking systematic exploration of the potential correlations between them. Moreover, current research on factors influencing bias mainly centers on how explicit bias changes, while the dynamics of implicit bias under these factors remain unexplored.
@@ 3 Methodology
LLMs possess the capability to reflect on and evaluate their own outputs Madaan et al. (2023); Weng et al. (2023); Shinn et al. (2023), offering a novel perspective for studying their biases. We leverage these reflective capabilities: first measuring the model’s implicit bias, then measuring its explicit bias by prompting the model to reflect on its implicit bias. This approach enables direct comparison between implicit and explicit biases toward identical social targets, thereby revealing the relation between these two types of bias within the same model. Specifically, our evaluation methodology draws inspiration from psychological measures including the Implicit Association Test (IAT) and Self-Report Assessment (SRA), which are used to measure implicit and explicit biases, respectively. Through prompt engineering, these methods are adapted for evaluating LLMs.
@@ 3.1 Implicit Bias Assessment Based on Psychological Theory
Research in social psychology suggests that the key to measuring implicit bias lies in conducting assessments without participants’ awareness Greenwald and Banaji (1995). The Implicit Association Test (IAT) is a major measurement tool developed on this principle: instead of directly asking participants about their attitudes toward social groups such as gender or race, it measures implicit bias by examining differences in response times when participants associate target-related stimuli (such as names of different genders or races) with corresponding attributes. For example, if a participant responds significantly faster to male-career and female-family pairings than to female-career and male-family pairings, this may indicate an implicit bias associating men with careers and women with family roles. To measure implicit bias within LLMs, we adapt the IAT to model evaluation through prompt engineering. Specifically, we propose a prompt template containing attribute words and masked targets:
<mask> is attrX as <mask> is attrY,
where $mask$ represents the masked target, while $attr_{X}$ and $attr_{Y}$ denote a predefined pair of attributes. This template requires the model to make a choice among given candidate words, thereby revealing its implicit bias. To illustrate this methodology, we use the implicit associations between racial categories (White-Black people) and valence attributes (good-bad) as an example. The experimental procedure consists of the following steps:
@@
    Define attribute word pairs $A_{X}$ and $A_{Y}$, representing good and bad respectively.
@@
    Randomly select five stimuli from $S_{a}$ and $S_{b}$ to form subsets $s_{a}$ and $s_{b}$, and randomly choose one attribute word each ($a_{X}$ and $a_{Y}$) from $A_{X}$ and $A_{Y}$.
@@
    Insert $a_{X}$ and $a_{Y}$ into the attribute positions of the template, then combine and shuffle $s_{a}$ and $s_{b}$ to create a candidate word set $s_{ab}$.
@@
    Design a prompt including the above elements with clear instructions, asking the model to select two words from $s_{ab}$ to fill the $mask$ positions in the template.
To enhance measurement reliability, we follow recommendations on prompt template variation from prior research Seshadri et al. (2022); Goldfarb-Tarrant et al. (2023) and employ several optimization techniques. First, we create five semantically similar but structurally distinct test templates through synonym substitution and word modification. These templates are provided in Appendix A. To further mitigate order effects, we swap the attribute pairs within each template, resulting in 10 templates. Beyond racial bias, we conduct comparative experiments on both explicit and implicit biases across multiple sensitive social dimensions including gender, occupation, age, and disability, to ensure comprehensive experimental coverage.
@@ 3.2 Explicit Bias Assessment Through Self-Reflection
To measure explicit bias in LLMs, we draw on Self-Report Assessment (SRA) from social psychology, a classic technique for measuring explicit bias. Unlike the indirect measurement of implicit bias, explicit bias measurement requires subjects to directly express their attitudes and views toward specific social groups while being aware of the measurement’s purpose. The Likert scale Likert (1932) is one of the most commonly used measurement tools in SRA, evaluating subjects’ attitudes with 5-point or 7-point rating scales.
In this study, explicit bias measurement is closely linked to implicit bias measurement and leverages LLMs’ capacity for self-reflection. Specifically, we measure explicit bias by guiding LLMs to systematically reflect on and evaluate their potential attitudes demonstrated during the implicit bias measurement phase. Since this process requires LLMs to analyze and evaluate their implicit biases shown in Section 3.1, we refer to this approach as the "self-reflection" assessment method.
During implementation, the measurement template explicitly specifies the target social groups, requiring the replacement of $mask$ in templates from the implicit test with specific social group concept words, such as men and women for gender categories, or Black and White for racial categories. Specifically, we randomly select words from the predefined social group word sets $T_{A}$ and $T_{B}$ to replace $mask$, while keeping the predefined attribute words $attribute_{X}$ and $attribute_{Y}$ unchanged, to form the following template sentence:
TargetA is attrX as TargetB is attrY.
Subsequently, the LLM is prompted to evaluate the template sentence using a 5-point Likert scale with options: strongly disagree, disagree, neutral, agree, strongly agree. The LLM assesses its level of agreement based on the specific content of the sentence, selects the most appropriate option from these five choices, and provides a reason for its selection. During the testing process, we randomize the five Likert scale options during each test rather than presenting them in a fixed order. This procedure helps reduce potential interference from option ordering on model choices, thereby maximizing the stability and consistency of measurement results.